OpenStudy (anonymous):

3x-15y=0 and 6x+10y=0. Are these lines parallel, intersecting, or one line?

OpenStudy (anonymous):

intersecting

OpenStudy (p0sitr0n):

x/5 =y -3/5x =y slope not same, so not parallel slope m1 not -1/m2m so not perpendicular obviously, not same line so, intersecting

OpenStudy (anonymous):

Did you determine that based on the fact they are both equal to zero? Or how did you get that?

OpenStudy (p0sitr0n):

no, in order to see this, isolate y and youll have a line of form y =mx +b here, b was 0 if lines are perpendicular m1 = -1/m2 if parallel, m1 =m2, but b1 not = b2 if same, same equation